# Can Masked Autoencoders Also Listen to Birds?

### v3

@@ Abstract
Masked Autoencoders (MAEs) have shown competitive results in audio classification by learning rich semantic representations through an efficient self-supervised reconstruction task. However, general-purpose models fail to generalize well when applied directly to fine-grained audio domains. Specifically, bird-sound classification requires distinguishing subtle inter-species differences and managing high intra-species acoustic variability, thereby revealing the performance limitations of general-domain Audio-MAE models. This work demonstrates that bridging this domain gap requires more than domain-specific pretraining data; adapting the entire training pipeline is crucial. We systematically revisit and adapt the pretraining recipe, fine-tuning methods, and frozen feature utilization to bird sounds using BirdSet, a large-scale bioacoustic dataset comparable to AudioSet. Our resulting Bird-MAE achieves new state-of-the-art results in BirdSet’s multi-label classification benchmark. Additionally, we introduce the parameter-efficient prototypical probing, enhancing the utility of frozen MAE representations and closely approaching fine-tuning performance in low-resource settings. Bird-MAE’s prototypical probes outperform linear probing by up to 37%${}_{\text{p}}$ in MAP and narrow the gap to fine-tuning to approximately 3.3%${}_{\text{p}}$ on average across BirdSet downstream tasks. Bird-MAE also demonstrates robust few-shot capabilities with prototypical probing in our newly established few-shot benchmark on BirdSet, highlighting the potential of tailored self-supervised learning pipelines for fine-grained audio domains.
@@ 1 Introduction
While fine-tuned Audio-MAEs demonstrate competitive performance on broad audio benchmarks beyond AudioSet, such as ESC-50 (Piczak, 2015) or Speech Commands (Warden, 2018), their direct transfer to fine-grained audio tasks is limited. For instance, general-purpose models exhibit a notable performance gap in specialized tasks such as bird sound classification compared to domain-specific supervised models (Ghani et al., 2023; Hamer et al., 2023) due to the diverging acoustic characteristics of bird calls versus general audio. Additionally, MAE’s reconstruction objective yields representations that require fine-tuning, providing limited utility as readily available frozen representations (Alkin et al., 2025), a drawback common among audio SSL methods (Chen et al., 2023; 2024a). Adapting a general-purpose model to new domains via fine-tuning is resource-intensive, requiring high computational resources and extensive labeled data (Han et al., 2024). Leveraging frozen representations with lightweight probes offers a more efficient solution, potentially achieving strong performance with less labeled data if the representations are truly domain-relevant. The distinct evolution of SSL for general audio versus speech processing (Baevski et al., 2020) further suggests benefits from a domain-adapted SSL model. Thus, achieving optimal performance with efficient models in fine-grained audio tasks necessitates more than off-the-shelf MAEs: it calls for a domain-aware adaptation of the entire SSL pipeline, including pretraining, fine-tuning, and frozen feature utilization.
Bird species classification exemplifies these challenges. It involves not only a domain shift from general audio but also the fine-grained task of distinguishing subtly different species with low inter-class variation amidst high intra-class variability (Rauch et al., 2024). These complexities are compounded by scarce labeled data (e.g., rare species) and computational constraints for edge deployment (Bellafkir et al., 2023). Current state-of-the-art (SOTA) models, such as Perch (Hamer et al., 2023), rely on supervised learning, failing to capitalize on the vast amounts of unlabeled bioacoustic data available for SSL. With the introduction of BirdSet (Rauch et al., 2024), a large-scale benchmark comparable to AudioSet but explicitly tailored to avian bioacoustics, bird sound classification emerges as an ideal scenario to investigate the benefits of a domain-specific SSL model. To address these challenges, this work examines the efficacy of a holistically adapted, domain-specific SSL pipeline for bird sound classification. We introduce the Bird-MAE model pretrained exclusively on bird vocalizations. Additionally, to better leverage its frozen representations in computationally efficient few-shot settings, we propose prototypical probing inspired by recent advancements in prototype-based learning (Heinrich et al., 2025). Our key contributions are summarized as follows:
@@
Contributions (1) We empirically demonstrate the domain gap between a general-purpose model and a domain-specific model. We emphasize the necessity of domain-specific SSL for fine-grained audio tasks. (2) We holistically adapt the MAE pipeline for bird sound classification. [^1] https://github.com/DBD-research-group/Bird-MAE We revisit design choices in the pretraining recipe, fine-tuning methods, and utilization of frozen representations. (3) We introduce and open-source Bird-MAE [^2] https://huggingface.co/DBD-research-group/Bird-MAE., a domain-specific MAE pretrained on BirdSet with various backbone sizes. Bird-MAE achieves SOTA results on BirdSet’s multi-label benchmark, improving MAP by an average of 15%${}_{\text{p}}$ across downstream tasks over prior best models. (4) We propose prototypical probing, a parameter-efficient probing technique that significantly enhances the utility of frozen MAE representations in bird sound classification. Prototypical probing boosts MAP by up to 37%${}_{\text{p}}$ over linear probing on BirdSet tasks, reducing the performance gap relative to full fine-tuning to an average of just 3.3%${}_{\text{p}}$. (5) We establish a novel few-shot multi-label classification benchmark for BirdSet. With prototypical probing, Bird-MAE enables performance in the few-shot benchmark close to that achieved using the full training dataset, highlighting its efficiency in low-data regimes.
@@ 2 Related Work
SSL in audio classification. SSL in audio classification has advanced the field, spanning from environmental sounds like ESC-50 (Piczak, 2015) to large-scale benchmarks like AudioSet that covers a wide range of sounds (e.g., human, animal, musical). Analogous to ImageNet (Deng et al., 2009) in vision, AudioSet provides a large-scale dataset for pretraining SSL models and evaluating their learned representations. While speech SSL models like Wav2vec2 (Baevski et al., 2020) usually operate on waveforms, general audio classification succeeds by adapting vision-based SSL techniques to spectrograms, achieving SOTA results on AudioSet (Chen et al., 2023; 2024a). MIM has successfully transitioned to audio classification by reconstructing masked spectrogram patches, introducing the Audio-MAE (Huang et al., 2022). This pretraining paradigm offers computational efficiency and fosters the learning of rich audio representations from unlabeled data. Subsequent work, such as BEATs (Chen et al., 2023) and EAT (Chen et al., 2024a), further progress audio MIM, incorporating a teacher-student approach. Our work centers on the MAE architecture. Its conceptual simplicity and pretraining efficiency make it an ideal baseline for isolating the impact of domain-specific adaptation and assessing the efficacy of novel probing techniques.
@@
Transfer learning in audio classification. General-purpose audio SSL models have proven effective for diverse downstream tasks (Turian et al., 2022; Saeed et al., 2021). BEATs (Chen et al., 2023), EAT (Chen et al., 2024a) and Audio-MAE (Huang et al., 2022) demonstrate their performance on tasks such as speech emotion recognition or environmental sound classification. However, recent studies reveal a notable performance degradation when these general-purpose models are applied to highly specialized domains, particularly bioacoustics (Hamer et al., 2023; Ghani et al., 2023). Benchmarks designed for transfer learning, such as HEAR (Turian et al., 2022), and bioacoustic benchmarks like BirdSet (Rauch et al., 2024) or BIRB (Hamer et al., 2023) highlight this limitation. Specifically, Audio-MAE (Huang et al., 2022) performs worse than spectrogram-based features from supervised models in bioacoustic tasks (Ghani et al., 2023). This performance drop underscores the current limitations of relying on general-domain pretraining for more fine-grained audio tasks and motivates the development of domain-specific solutions. In this work, we address this limitation by introducing and evaluating Bird-MAE specifically adapted to bird sound classification, quantifying the benefits of domain-specific solutions for fine-grained classification in audio.
@@
Downstream task adaptation in SSL. Adapting models to downstream tasks typically involves full fine-tuning or utilizing frozen representations with lightweight probes (Marks et al., 2025). While fine-tuning often yields the highest performance, it can be computationally expensive and may lead to overfitting on smaller datasets. Thus, utilizing frozen representations has gained notable interest in vision (Oquab et al., 2024; El-Nouby et al., 2024; Xie et al., 2022; Assran et al., 2023). Standard approaches involve extracting features (e.g., via global average pooling or the cls-token) and training a simple classifier, such as linear probing (Oquab et al., 2024), k-NN probing (Zhou et al., 2021; Kakogeorgiou et al., 2022; Lehner et al., 2024), or shallow MLP probing (Dubois et al., 2022; Fuller et al., 2023; Tschannen et al., 2023). However, it is widely observed that representations learned via generative tasks like MIM underperform with linear probing compared to contrastive methods (Alkin et al., 2025; He et al., 2022; Park et al., 2023). To mitigate this gap in MIM, upstream feature refinement (Alkin et al., 2025) or alternative probing methods have been explored in vision. Notably, attentive probing (El-Nouby et al., 2024; Lee et al., 2019) applies an attention mechanism over patch tokens and improves frozen representations with low computational overhead (Yu et al., 2022; Chen et al., 2024b; Darcet et al., 2025). Another probing paradigm involves prototypical networks (Snell et al., 2017; Palanisamy et al., 2024; Tian et al., 2024), which extract class centroids from frozen representations for a similarity-based class assignment without retraining. Despite these advancements in vision, current best-performing audio SSL models (Huang et al., 2022; Chen et al., 2023; 2024a) rely on full-model fine-tuning, suggesting their frozen representations offer suboptimal performance or are not properly utilized. In our setting, we evaluate standard parametric probing methods for downstream adaptation using frozen representations, including linear, MLP, and attentive probing. Additionally, we propose and analyze prototypical probing, adapted from prototypical networks in bioacoustics (Heinrich et al., 2025), designed to utilize spatial MAE features effectively.
@@
Bird sound classification. Supervised learning has dominated research in bird sound classification, typically employing convolutional architectures that remain the top performers on bioacoustic benchmarks (Stowell, 2021; Rauch et al., 2024). For instance, Google’s Perch model (Hamer et al., 2023) is based on the EfficientNet architecture (Tan & Le, 2019). The feasibility of large-scale supervised training stems from community-driven platforms like Xeno-Canto (XC) (Vellinga & Planqué, 2015), which currently hosts over 850k weakly-labeled bird recordings. Perch and BirdNeXt (Rauch et al., 2024) derive their training data from XC. However, reliance on manually curated, non-standardized datasets from these platforms has hindered comparison across studies and methods (Rauch et al., 2024). The introduction of the BirdSet dataset and multi-label bird classification benchmark, which contains volumes of pretraining data comparable to AudioSet, makes this comparison possible for domain-specific SSL. While SSL has shown promise in speech and general audio classification, its evaluation in bioacoustics is less mature. NatureLM-audio (Robinson et al., 2025) presents the first audio-language model tailored to general bioacoustics, demonstrating competitive zero-shot performance on BirdSet. Existing domain-specific SSL models for bioacoustics, such as BirdAVES (Hagiwara, 2023) and contrastive models (Moummad et al., 2024), have not yet been evaluated under the standardized conditions provided by BirdSet. This work introduces the domain-specific Bird-MAE and comprehensively evaluates it on BirdSet, establishing a new SOTA.
@@ 3.1 Pretraining (M1)
Pretraining lays the foundation for effective SSL by learning representations adaptable to downstream tasks. The core of the Audio-MAE baseline is the pretrained encoder ${h}_{\alpha}:\mathcal{X}\subseteq\mathbb{R}^{F\times T}\rightarrow\mathbb{R}^{H\times W\times D}$ with parameters $\alpha$ from AudioSet. Here, $F$ represents the number of frequency bins and $T$ the number of time frames of an input spectrogram $\mathbf{x}\in\mathcal{X}$. The encoder maps $\mathbf{x}$ to a patch-based feature map $\mathbf{h}_{\alpha}(\mathbf{x})$, where $H$ and $W$ denote the number of non-overlapping patches along the height and width dimensions, and $D$ is the feature dimension per patch. For instance, given an AudioSet spectrogram image with $F\!=\!128$ and $T\!=\!1024$ with a patch size of $16\!\times\!16$, the encoder $h_{\alpha}$ produces a feature map $\mathbf{h}_{\alpha}(\mathbf{x})$ of dimension $8\!\times\!64\!\times\!D$. Our proposed modifications result in an encoder $h$, pretrained on bird sounds, denoted ${h}_{\beta}$. The key changes from the baseline Audio-MAE pretraining from Huang et al. (2022) are listed in Table 1 with the detailed ablations in Section 5.2. The modifications within this module include:
Model Dataset Img Decoder Epochs Masking Batch Mean Std Mixup LR WD
Audio-MAE AS-2M 1024$\times$128 Swin 32 0.8 512 -4.2 4.569 0 2e-4 1e-4
Bird-MAE XCL-1.7M 512$\times$128 ViT 150 0.75 1024 -7.2 4.43 0.3 2e-4 1e-4
@@
Data source. The choice of pretraining data influences downstream performance, especially when adapting models from coarse-grained to fine-grained classification tasks. General-purpose datasets like AudioSet encompass a broad spectrum of acoustically distinct classes (high inter-class variation), encouraging models to learn discriminative general features. However, fine-grained domains like bird sound classification require distinguishing between subtle different species (low inter-class variation) while handling acoustic variability within each species (high intra-class variation) (Rauch et al., 2024). AudioSet, despite including some animal sounds, does not adequately prepare models for these fine-grained bioacoustic nuances. Therefore, to develop a model tailored for this challenge, we replace AudioSet with domain-specific pretraining data derived from BirdSet (XCL-1.7M after curation, see Section 4).

### v4

@@ Abstract
Masked Autoencoders (MAEs) learn rich semantic representations in audio classification through an efficient self-supervised reconstruction task. However, general-purpose models fail to generalize well when applied directly to fine-grained audio domains. Specifically, bird-sound classification requires distinguishing subtle inter-species differences and managing high intra-species acoustic variability, revealing the performance limitations of general-domain Audio-MAEs. This work demonstrates that bridging this domain gap domain gap requires full-pipeline adaptation, not just domain-specific pretraining data. We systematically revisit and adapt the pretraining recipe, fine-tuning methods, and frozen feature utilization to bird sounds using BirdSet, a large-scale bioacoustic dataset comparable to AudioSet. Our resulting Bird-MAE achieves new state-of-the-art results in BirdSet’s multi-label classification benchmark. Additionally, we introduce the parameter-efficient prototypical probing, enhancing the utility of frozen MAE representations and closely approaching fine-tuning performance in low-resource settings. Bird-MAE’s prototypical probes outperform linear probing by up to 37 percentage points in mean average precision and narrow the gap to fine-tuning across BirdSet downstream tasks. Bird-MAE also demonstrates robust few-shot capabilities with prototypical probing in our newly established few-shot benchmark on BirdSet, highlighting the potential of tailored self-supervised learning pipelines for fine-grained audio domains.
@@ 1 Introduction
While fine-tuned Audio-MAEs demonstrate competitive performance on audio benchmarks beyond AudioSet, such as ESC-50 (Piczak, 2015), their direct transfer to more fine-grained audio tasks is limited. For instance, general-purpose models exhibit a notable performance gap in specialized tasks such as bird sound classification compared to domain-specific supervised models (Ghani et al., 2023; Hamer et al., 2023). Although AudioSet also contains bird sounds, its coarse-grained nature fails to equip pretrained models with the fine-grained discriminative information required for bird sound classification. In this domain, models must distinguishing between acoustically similar species (low inter-class variation) while handling diverse vocalizations within a single species (high intra-class variation) (Rauch et al., 2024b). Compounding this domain mismatch, MAE’s reconstruction objective yields representations that require fine-tuning, offering limited utility as frozen features (Alkin et al., 2025), a drawback common in audio SSL (Chen et al., 2023; 2024a). This presents a critical trade-off: Adapting a general-purpose model via full fine-tuning on different tasks with scarcely labeled data is resource-intensive (Han et al., 2024), while creating a domain-specific foundation model from scratch entails an upfront computational investment. For domains where general audio models seem to plateau (Turian et al., 2022), this one-time investment is justified by the benefits unlocked across the downstream lifecycle. A domain-specific foundation model enables efficient adaptation to various sub-tasks (e.g., population density or call-type classification in bioacoustics) via lightweight probing of frozen representations (Ghani et al., 2023). This efficiency is crucial in fields like bioacoustics, where researchers (e.g., biologists) often face scarce labels and have limited computational resources for edge deployment (Bellafkir et al., 2023).
Given these challenges of fine-grained classification and the need for downstream efficiency, bird sound classification emerges as an ideal testbed for investigating our core hypothesis: achieving state-of-the-art (SOTA) performance in audio currently requires a holistic, domain-aware adaptation of the entire SSL pipeline, including pretraining, fine-tuning and frozen feature utilization. Current SOTA models in this area, such as Perch (Hamer et al., 2023), still rely on supervised learning, failing to capitalize on the vast unlabeled bioacoustic data suitable for SSL. The large-scale BirdSet benchmark (Rauch et al., 2024b), comparable pretraining data volumes to AudioSet, provides the necessary resources to test this hypothesis. Thus, we introduce Bird-MAE, a model pretrained exclusively on bird vocalizations via a fully adapted SSL pipeline. To better leverage its frozen representations, we propose a novel application of prototype learning (Heinrich et al., 2025) as an efficient probing mechanism (i.e., prototypical probing) and systematically evaluate it in low- and high-data regimes. Our key contributions are summarized as follows:
@@
Contributions (1) We empirically demonstrate the domain gap between a general-purpose model and a domain-specific model. We emphasize the necessity of domain-specific SSL for fine-grained audio tasks. (2) We holistically adapt the MAE pipeline for bird sound classification. [^1] https://github.com/DBD-research-group/Bird-MAE We revisit design choices in the pretraining recipe, fine-tuning methods, and utilization of frozen representations. (3) We introduce Bird-MAE [^2] https://huggingface.co/collections/DBD-research-group/bird-mae-68a4654a4cc4d363eb87e6f2., a domain-specific MAE pretrained on BirdSet. Bird-MAE achieves SOTA results on BirdSet’s multi-label benchmark, improving mean avarage precision (mAP) by an average of 15 percentage points ($p$) across downstream tasks over prior best models. (4) We introduce prototypical probing, an application of prototype-based layers as a parameter-efficient technique for probing frozen MAE features. We systematically benchmark it against linear, MLP and attentive probes, including few-shot settings. This method lifts mAP by up to 37 pp over linear probing and leaves only a 3.3 $p$ performance gap to full fine-tuning across BirdSet tasks. (5) We establish a novel few-shot multi-label classification benchmark for BirdSet. With prototypical probing, Bird-MAE enables performance in the few-shot benchmark close to that achieved using the full training dataset, highlighting its efficiency in low-data regimes.
@@ 2 Related Work
SSL in audio classification. SSL in audio classification has advanced the field, spanning from environmental sounds like ESC-50 (Piczak, 2015) to large-scale benchmarks like AudioSet that covers a wide range of sounds (e.g., human, animal, musical). Analogous to ImageNet (Deng et al., 2009) in vision, AudioSet provides a large-scale dataset for pretraining SSL models and evaluating their learned representations. While speech SSL models like Wav2vec2 (Baevski et al., 2020) usually operate on waveforms, general audio classification succeeds by adapting vision-based SSL techniques to spectrograms, achieving SOTA results on AudioSet (Chen et al., 2023; 2024a). MIM has successfully transitioned to audio classification by reconstructing masked spectrogram patches, introducing the Audio-MAE (Huang et al., 2022). This pretraining paradigm offers computational efficiency and fosters the learning of rich audio representations from unlabeled data. Subsequent work, such as BEATs (Chen et al., 2023) and EAT (Chen et al., 2024a), further progress audio MIM, incorporating a teacher-student approach. Our work centers on the MAE architecture. Its conceptual simplicity and pretraining efficiency make it an ideal baseline for isolating the impact of domain-specific adaptation and assessing the efficacy of novel probing techniques in challenging fine-grained domains.
@@
Transfer learning in audio classification. General-purpose audio SSL models have proven effective for diverse downstream tasks (Turian et al., 2022; Saeed et al., 2021). BEATs (Chen et al., 2023), EAT (Chen et al., 2024a) and Audio-MAE (Huang et al., 2022) demonstrate their performance on tasks such as speech emotion recognition or environmental sound classification. However, recent studies reveal a notable performance degradation when these general-purpose models are applied to highly specialized domains, particularly bioacoustics (Hamer et al., 2023; Ghani et al., 2023). Benchmarks designed for transfer learning, such as HEAR (Turian et al., 2022), and bioacoustic benchmarks like BirdSet (Rauch et al., 2024b) or BIRB (Hamer et al., 2023) highlight this limitation. Specifically, Audio-MAE (Huang et al., 2022) performs worse than spectrogram-based features from supervised models in bioacoustic tasks (Ghani et al., 2023). This performance drop underscores the current limitations of relying on general-domain pretraining for more fine-grained audio tasks and motivates the development of domain-specific solutions. In this work, we address this limitation by introducing and evaluating Bird-MAE specifically adapted to bird sound classification, quantifying the benefits of domain-specific solutions for fine-grained classification in audio.
@@
Downstream task adaptation in SSL. Adapting models to downstream tasks typically involves full fine-tuning or utilizing frozen representations with lightweight probes (Marks et al., 2025). While fine-tuning often yields the highest performance, it can be computationally expensive and may lead to overfitting on smaller datasets. Thus, utilizing frozen representations has gained notable interest in vision (Oquab et al., 2024; El-Nouby et al., 2024; Xie et al., 2022; Assran et al., 2023). Standard approaches involve extracting features (e.g., via global average pooling or the cls-token) and training a simple classifier, such as linear probing (Oquab et al., 2024), k-NN probing (Zhou et al., 2021; Kakogeorgiou et al., 2022; Lehner et al., 2024), or shallow MLP probing (Dubois et al., 2022; Fuller et al., 2023; Tschannen et al., 2023). However, it is widely observed that representations learned via generative tasks like MIM underperform with linear probing compared to contrastive methods (Alkin et al., 2025; He et al., 2022; Park et al., 2023). To mitigate this gap in MIM, upstream feature refinement (Alkin et al., 2025) or alternative probing methods have been explored in vision. Notably, attentive probing (El-Nouby et al., 2024; Lee et al., 2019) applies an attention mechanism over patch tokens and improves frozen representations with low computational overhead (Yu et al., 2022; Chen et al., 2024b; Darcet et al., 2025). Another probing paradigm involves prototypical networks (Snell et al., 2017; Palanisamy et al., 2024; Tian et al., 2024), which extract class centroids from frozen representations for a similarity-based class assignment without retraining. Despite these advancements in vision, current best-performing audio SSL models (Huang et al., 2022; Chen et al., 2023; 2024a) rely on full-model fine-tuning, suggesting their frozen representations offer suboptimal performance or are not properly utilized. In our setting, we evaluate standard parametric probing methods for downstream adaptation using frozen representations, including linear, MLP, and attentive probing. Additionally, we propose and analyze prototypical probing, utilized from prototypical networks in bioacoustics (Heinrich et al., 2025) for a new purpose: as a lightweight, parameter-efficient probe for frozen MAE representations, effectively utilizing their spatial features.
@@
Bird sound classification. Supervised learning has dominated research in bird sound classification, typically employing convolutional architectures that remain the top performers on bioacoustic benchmarks (Stowell, 2021; Rauch et al., 2024b). For instance, Google’s Perch model (Hamer et al., 2023) is based on the EfficientNet architecture (Tan & Le, 2019). The feasibility of large-scale supervised training stems from community-driven platforms like Xeno-Canto (XC) (Vellinga & Planqué, 2015), which currently hosts over 850k weakly-labeled bird recordings. Perch and BirdNeXt (Rauch et al., 2024b) derive their training data from XC. However, reliance on manually curated, non-standardized datasets from these platforms has hindered comparison across studies and methods (Rauch et al., 2024b). The introduction of the BirdSet dataset and multi-label bird classification benchmark, which contains volumes of pretraining data comparable to AudioSet, makes this comparison possible for domain-specific SSL. While SSL has shown promise in speech and general audio classification, its evaluation in bioacoustics is less mature. NatureLM-audio (Robinson et al., 2025) presents the first audio-language model tailored to general bioacoustics, demonstrating competitive zero-shot performance on BirdSet. Existing domain-specific SSL models for bioacoustics, such as BirdAVES (Hagiwara, 2023) and contrastive models (Moummad et al., 2024), have not yet been evaluated under the standardized conditions provided by BirdSet. This work introduces the domain-specific Bird-MAE and comprehensively evaluates it on BirdSet, establishing a new SOTA.
@@ 3.1 Pretraining (M1)
Pretraining lays the foundation for effective SSL by learning representations adaptable to downstream tasks. The core of the Audio-MAE baseline is the pretrained encoder ${h}_{\alpha}:\mathcal{X}\subseteq\mathbb{R}^{F\times T}\rightarrow\mathbb{R}^{H\times W\times D}$ with parameters $\alpha$ from AudioSet. Here, $F$ represents the number of frequency bins and $T$ the number of time frames of an input spectrogram $\mathbf{x}\in\mathcal{X}$. The encoder mAPs $\mathbf{x}$ to a patch-based feature map $\mathbf{h}_{\alpha}(\mathbf{x})$, where $H$ and $W$ denote the number of non-overlapping patches along the height and width dimensions, and $D$ is the feature dimension per patch. For instance, given an AudioSet spectrogram image with F=128F\!=\!128 and T=1024T\!=\!1024 with a patch size of $16\!\times\!16$, the encoder $h_{\alpha}$ produces a feature map $\mathbf{h}_{\alpha}(\mathbf{x})$ of dimension $8\!\times\!64\!\times\!D$. Our proposed modifications result in an encoder $h$, pretrained on bird sounds, denoted ${h}_{\beta}$. The key changes from the baseline Audio-MAE pretraining from Huang et al. (2022) are listed in Table 1 with the detailed ablations in Section 5.2. The modifications within this module include:
[…]
@@
Data source. The choice of pretraining data influences downstream performance, especially when adapting models from coarse-grained to fine-grained classification tasks. General-purpose datasets like AudioSet encompass a broad spectrum of acoustically distinct classes (high inter-class variation), encouraging models to learn discriminative general features. However, fine-grained domains like bird sound classification require distinguishing between subtle different species (low inter-class variation) while handling acoustic variability within each species (high intra-class variation) (Rauch et al., 2024b). AudioSet, despite including some animal sounds, does not adequately prepare models for these fine-grained bioacoustic nuances. Therefore, to develop a model tailored for this challenge, we replace AudioSet with domain-specific pretraining data derived from BirdSet (XCL-1.7M after curation, see Section 4).